OpenStudy (anonymous):

y-8=-1.5(x+4) in standard form (in point slope form already)

OpenStudy (anonymous):

ok well first distribute the 1.5 with everything in the parentheses

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

y-8=-1.5(x+4) y-8=-1.5x-6 ok now add 8 to the -6

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok so now ur equation is y=-1.5x+2 your slope is -1.5 and your y intercept is 2 do u need to graph them?

OpenStudy (anonymous):

no i need it in standard form as in Ax + By = C

OpenStudy (anonymous):

ooohh so add 1.5x to the other side do you know what it would be then?

OpenStudy (anonymous):

nope.... sorry not good at math

OpenStudy (anonymous):

its ok want me to tell you?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

1.5x+1y=2 did you see how i got that?

OpenStudy (anonymous):

ooooooooh okay thanks! that was my last problem i was pretty stumped lol

OpenStudy (anonymous):

haha. but do u see how I got that?

OpenStudy (anonymous):

yes bye thank you

OpenStudy (anonymous):

welcome